### v1

# Subject-Oriented Video Captioning

Abstract
Describing video content according to users’ needs is a long-held goal. Although existing video captioning methods have made significant progress, the generated captions may not focus on the entity that users are particularly interested in. To address this problem, we propose a new video captioning task, subject-oriented video captioning, which allows users to specify the describing target via a bounding box. To support this task, we construct two subject-oriented video captioning datasets based on two widely used video captioning datasets: MSVD and MSRVTT, by annotating subjects in each video for each caption. These datasets pave the way for future technique development. As the first attempt, we evaluate four state-of-the-art general video captioning models, and have observed a large performance drop. We then explore several strategies to enable them to describe the desired target. Experimental results show obvious improvement, but there is still a large room for further exploration in this field.
1 Introduction
Figure 1: Conventional video captioning (upper panel) vs. our subject-oriented video captioning (bottom panel).
Video captioning is an important video understanding task, aiming to accurately describe the video content in a single sentence. It plays a crucial role in numerous applications, such as video title generation [35], blind assistance [6], and video search [13]. However, existing methods do not allow users to specify the target of interest so as to generate desired captions. This could largely limit its application in real-world scenarios. Consider a video where multiple individuals are present as shown in Figure 1, if a user needs a summary of the behaviors of a specific individual, such as the woman, traditional video captioning methods have a large chance to fail, e.g., the state-of-the-art methods SwinBERT [9] and HMN [32] generate “A man is talking about a car” and “A man is driving a car”, respectively. This is because they are trained to cover the majority of entities mentioned in the ground truth, which is “a man” in this case. Consequently, they cannot cater to user-specific needs for targeted descriptions.
In the field of image captioning, similar attempts have been made. For instance, a framework [4] named “Show, Control and Tel” generates captions by taking a series of image regions as input. However, this method requires users to select and rank multiple entity areas, which is impractical in the video scenario because it is unreasonable to require the user to watch the whole video first and then pick out interacted objects. A desired method should just require to specify the target of interest and let the model its own to discover what happens to the target. Another limitation of applying image captioning methods for video captioning is their inability to handle temporal information in multiple frames, such as complex actions. The above mentioned limitations also exist in other image captioning methods, such as CGO [37] and VSR-guided model [3].
Another line of related research is controlled video captioning. Existing methods primarily focus on controlling sentence structure, as demonstrated by [25], which introduces Part-of-Speech(POS) information for this purpose. SCVC [34] aims to generate one caption that describes the video contents semantically while also imitating the syntactic structure of the given example sentence. However, these methods lack effective control over the subjects in the videos. Meanwhile, although O2NA [11] allows for control using simple entity-class words, it does not enable the selection of specific entity areas, thus lacking flexibility and failing to meet our requirements.
To address the aforementioned problems, we propose a new task named “subject-oriented video captioning”, which aims to generate descriptions centered on activities of the user-specified entity. One key reason limiting the development of such desired methods is the lack of proper data. Existing video captioning datasets do not provide video annotations corresponding to the subject in each caption. This raises difficulty in learning the correspondence between captions and numerous entities in a video. In this work, we annotate two widely used video captioning datasets: MSVD [2] and MSR-VTT [26]. This paves the way to develop methods that generate captions describing user-specified targets. As an example, in the bottom panel of Figure 1, trained with the new data, our method can generate descriptions closely related to each specified entity. For example, when specifying the car as the description target, the generated caption is “A car is being shown”; when the target of interest is changed to the woman, the generated caption changes to “A woman is driving a car”.
To verify the necessity of developing new methods in this direction, we evaluate four state-of-the-art conventional video captioning methods and observe significant performance degradation on the new task: 20$\sim$30 absolute drop in terms of the main metric Cider on our subject-oriented MSVD, and 6$\sim$51 absolute drop on our subject-oriented MSRVTT. This shows there is ample room for exploration in this field. As our initial endeavor, we have undertaken several preliminary attempts where we integrated subject features into existing frameworks based on Transformer and LSTM. These preliminary attempts have yielded certain successes and may serve as inspiration for future research.
The main contributions of this paper are as follows:
  • We propose a new task, subject-oriented video captioning, and build two new datasets based on MSVD and MSR-VTT, which paves the way for developing more flexible video captioning methods.
  • We evaluate four representative conventional video captioning methods, and observe large performance degradation, which shows the necessity of developing new techniques.
  • We carry out extensive preliminary attempts to address the new task and achieve favorable improvements, which may serve as a strong baseline and inspire future research.
Figure 2: Conventional video captioning dataset (left panel) vs. our Subject-oriented video captioning dataset (right panel).
2 Related Work
In this section, we briefly review general video captioning methods and controllable video captioning methods, which are related to our task and put this work in a proper context.
2.1 General Video Captioning
Recent research on video captioning primarily focuses on the design of network architectures. Early methods employ simple CNN architectures for representation learning from videos, such as S2VT [24]. Later, more complex encoder architectures are proposed. STG-KD [14] and ORG-TRL [36] design encoders based on object relation graphs to distinguish the spatial-temporal relationships among video objects. SGN [19] proposes a semantic grouping network, which encodes videos into semantic clusters based on relevant frames. HMN [32] adopts a hierarchical modular design to encode three levels of visual features to get comprehensive video representations. GL-RG [27] creates a global-local encoder, which exploits rich temporal representation for video captioning. Some work also focuses on improving the decoding ability, such as the integration of temporal attention mechanisms by Yap et al. [31] and attended memory decoders by Pei et al. [16], marking significant enhancements in captioning quality. In addition to obtaining richer visual representations and enhancing decoding capabilities, the long-tailed problem hinders the effectiveness of low-frequency tokens, which play a crucial role in generating detailed content. RSFD [38] is proposed to constantly capture the linguistic representation of the infrequent tokens. TextKG [5] integrates an additional knowledge graph to mitigate the challenge of long-tail words.
Different from the aforementioned methods, SwinBERT [9] employs an end-to-end transformer-based design, which directly takes a sequence of video frames as inputs and uses caption-related loss to adjust feature learning. However, it increases the computational burden and time consuming to handle the dense video tokens. To enhance inference efficiency, Cocap [21] directly takes compressed video as input and fast generates captions. Most recently, some work [7, 20] leverages large pre-train models as the backbone to align powerful vision and language representations.
Although conventional video captioning methods, as mentioned above, are capable of generating accurate overall descriptions of a video clip, they fall short when it comes to emphasizing a particular entity of interest as the subject. In this work, we pioneer the task of subject-oriented video captioning, shifting the focus towards the activities of a specific subject within a video, rather than pursuing depicting general information in the video. Moreover, to accommodate our novel task, we have restructured the traditional captioning datasets and explore several attempts based on state-of-the-art general video captioning methods.
2.2 Controllable Video Captioning
In contrast to general video captioning, much less attention is paid to controllable video captioning. Some efforts involve using additional sentences as reference signals, allowing the generated sentences to learn the syntactic structure or style of the reference sentences. For example, SMCG [33] models the syntactic structure from a given exemplar sentence and designs a modified LSTM module to transfer the syntax in the exemplar to new captions for the input video. FS-StyleCap [29] extract the style of stylized example to generate sentences with the same style. On the other hand, some works employ special inputs as control signals. MSG [30] uses a controllable masked scene graph to represent interested objects and develops a corresponding decoder to infer masked relationships between objects to obtain the final caption. In [11], a method to generate captions about the specified objects is developed. To this end, object detection is performed, and then a non-autoregressive generator has to be used to make up a complete sentence from ahead-determined objects. ERA [10] introduces the reference videos to capture contextual information and further refine the group video captioning.
Our task/method differs from the above-mentioned methods in two aspects: (I) The target selection is not limited to the nodes of a scene graph of MSG. Our task/method allows the user to select any entity in the video as the desired caption subject. Furthermore, the caption content is from the whole video, not limited to all the graph nodes of MSG. (II) The additional control signals we utilized are unique, setting our approach apart from previous methods.
3 Subject-Oriented Video Captioning Task
In this paper, we introduce a new task, termed as “Subject-Oriented Video Captioning”. Unlike conventional video captioning tasks that generate a broad description of a video, our task aims to generate captions that predominantly focus on the activities of a specific entity specified by the user using a bounding box. To adapt to this newly devised task, traditional video captioning datasets have been meticulously refined, annotated, and reorganized, ensuring that they are well-suited to facilitate and nurture the generation of subject-oriented video captions. This enhances the specificity and relevance of the textual descriptions, echoing the nuanced activities of the identified subjects in videos.
In the next section, we provide the details of the dataset construction and statistics. At last, we evaluate four state-of-the-art conventional video captioning methods and our extensions on top of them.
Dataset Train Val Test Region Frame Caption
SO-MSVD 1,871 150 1,287 4,287 3,411 72,530
SO-MSRVTT 15,306 1,085 6,979 35,499 25,988 149527
Table 1: Statistics of our Subject-Oriented MSVD and MSRVTT datasets.
4 Dataset
4.1 Dataset Construction
To facilitate future research on the Subject-Oriented Video Captioning task, we construct datasets based on two widely used general video captioning datasets: MSVD [2] and MSRVTT [26].
We annotate the dataset in four steps. First, we use spaCy [^1] https://spacy.io/ to extract entities in each caption. We term these extracted entities as subjects. Captions with abstract subjects like “a video” are discarded. Then, we annotate the bounding box in videos for each subject. We only annotate the frame where the subject appears for the first time. To facilitate the annotation, we perform object detection using Faster-RCNN, then calculate and rank the cosine similarity between the entity categories predicted by Faster-RCNN and the subject words. Next, we manually check each bounding box for each caption. Missing and incorrect bounding boxes are corrected during this step. Finally, we reorganize the dataset as $\{V,S,C\}$, where $S$ represents the annotated subject regions within videos, $V$ denotes the video, and $C$ is the captions corresponding to the subject. As shown in right panel of Figure 2, one video can have multiple subjects, and each subject may have multiple captions.
4.2 Dataset Statistics
Table 1 presents the detailed specifics of the reannotated and reorganized datasets. In the subsequent sections, we elaborate on each of them, drawing upon the data shown in the table.
Subject-Oriented MSVD After filtering out captions with abstract subjects from the original MSVD dataset [2], we annotate 3,308 subject samples for 1,961 videos, with 72,530 captions. On average, one subject corresponds to 22 sentences. As shown in Figure 3, subject-oriented MSVD contains a varying number of subjects ranging from 1 to 9, with the corresponding average number of captions approximately between 3 and 38. Note that one subject may correspond to multiple frames and regions in a video. For example, consider the scenario where two men are talking about something as shown in Figure 4; these ‘two men’ may first appear in different frames, and each may occupy different regions even in the same frame. Based on our statistics, a total of 4,287 regions have been annotated across 3,411 frames. Since the total number of videos remains similar to its original scale, we adopt the same dataset division, leading to 1871, 150, and 1287 subject samples in the training, validation, and test splits, respectively. We also provide word-cloud-based statistics in Figure 5 (a) to reveal the relative amount of different subjects. It shows that the top-4 subjects in MSVD are “man”, “girl”, “person”, and “cat”, followed by “boy”, “dog”, “men”, and “monkey”.
Figure 3: Average amount of captions regarding subject complexity.
Figure 4: Multiple entities appear in different frames
Subject-Oriented MSRVTT We apply the same annotation strategy to the newly built training, validation, and test splits contain 15,306, 1,085, and 6,979 subject samples, respectively, from 6,310, 487, and 2,927 videos, respectively. Figure 3 illustrates the range of subject numbers in each video of subject-oriented MSRVTT, varying from 1 to 10, with the corresponding average number of captions approximately between 1 and 15. On this dataset, a total of 35,499 regions have been annotated across 25,988 frames. Figure 5 (b) presents the relative amount of different subjects in the reconstructed MSRVTT dataset. It shows that the top-4 subjects are slightly different from those of the new MSVD, where ‘people’ substitutes ‘cat’. We also observe that ‘men’, ‘chef’, ‘car’, ’player’, ’guy’, ‘women’, ‘children’, and ‘character’ have a similar amount, just next to the top-4 subjects.
(a) Subject-Oriented MSVD
(b) Subject-Oriented MSRVTT
Figure 5: Word cloud of our Subject-Oriented datasets. The bigger the font, the more percentage it occupies.
5 Method
Figure 6: The top row shows the original method, while the bottom row displays our improved approach.
To test the difficulty of the subject-oriented video captioning task, we choose four representative state-of-the-art video captioning methods, including AR-B [28], HMN [32], GL-RG [27] and SwinBERT [9]. According to the decoder architecture, HMN and GL-RG both adopt LSTM-based frameworks, whereas AR-B and SwinBERT are representative approaches utilizing Transformers [22]. Moreover, we also extended these four baselines to make them more suitable to address the newly proposed task.
In the following sections, we first briefly introduce the four baselines in Sec. 5.1. Then, we describe the specific adaptation of LSTM-decoder-based models to our task in Sec 5.2. Last, we present the modification of transformer-based methods in Section 5.3.
5.1 Baseline Methods
AR-B [28] belongs to traditional autoregressive captioning models, which use offline ResNet-101 image features and C3D motion features as inputs and employ a transformer decoder as a text generator.
GL-RG [27] proposes a precisely formulated global-local encoder to exploit rich temporal representation for video captioning.
HMN [32] proposes a hierarchical modular network that serves as a strong video encoder, which reduces the gap between video and language.
SwinBERT [9] is the first end-to-end fully Transformer-based model for video captioning. Different from previous methods leveraging off-the-shelf 2D appearance features and 3D motion features, it adopts a trainable VidSwin [12] as a feature extractor to encode spatial-temporal representations.
5.2 LSTM-based Subject-Oriented Methods
The structure of the LSTM-based method is shown in Figure 6 (a). , HMN and GL-RG belong to this category. They both consist of a Video Encoder and an LSTM Decoder to process and generate the video captions, respectively. The caption generation process is as follows:
[…]
$F_{v}=E_{v}(V).$ (1)
The structure of the encoder varies across methods. For instance, HMN employs a complex encoder comprising both Transformer and BiLSTM. For ease of representation, we uniformly denote as $E_{v}$.
Then, video features $F_{v}$ and the word $x_{t-1}$ generated at the previous time step are fed into an ${{\rm LSTM}}$ to predict word $x_{t}$ for the current time step:
${}x_{t}={{\rm LSTM}}(F_{v},x_{t-1}).$ (2)
[…]
$F_{s}=E_{s}(I).$ (3)
The encoder’s structure includes a pre-trained feature extractor and a linear layer for dimensionality reduction. For ease of description, we uniformly denote this entire structure as $E_{s}$. Then, we concatenate the encoded subject features with the original encoded video features
$F_{m}=Concat(F_{v},F_{s}),$ (4)
which is fed into the LSTM decoder Eq. 2 for the next word prediction.
Models Year SO-MSVD SO-MSRVTT
B@4 M R C B@4 M R C
AR-B [28]_ori_perf 2021 48.7 33.3 - 91.8 40.5 28.7 - 49.1
AR-B 2023 27.4 22.4 50.5 69.7 17.0 16.6 38.7 43.3
AR-B_ext 2023 33.5 25.3 53.8 86.9 19.6 17.8 40.6 50.2
HMN [32]_ori_perf 2022 59.2 37.7 75.1 104.0 43.5 29.0 62.7 51.5
HMN 2023 29.8 22.9 51.2 77.7 17.1 16.5 39.1 44.9
HMN_ext 2023 34.3 24.8 54.3 88.1 20.2 17.9 42.8 53.1
GL-RG [27]_ori_perf 2022 45.5 30.1 62.6 94.3 55.5 37.8 74.7 51.2
GL-RG 2023 34.5 26.1 54.4 86.7 18.3 17.3 39.1 43.8
GL-RG_ext 2023 37.2 27.8 56.9 97.6 22.2 18.9 43.9 58.6
SwinBert [9]_ori_perf 2022 55.7 39.7 75.7 109.4 41.9 29.9 62.1 53.8
SwinBert 2023 33.4 25.7 54.2 89.6 18.2 17.4 40.4 46.9
SwinBert_ext 2023 36.1 27.2 56.3 98.9 21.5 19.1 44.0 58.8
Table 2: Results on our Subject-Oriented MSVD and MSR-VTT datasets. Rows with grey background denote results on original MSVD and MSRVTT datasets. The best results are highlighted in bold.
5.3 Transformer-based Subject-Oriented Methods
The video captioning methods based on the Transformer architecture can be further divided into two categories: Transformer encoder-based methods and Transformer decoder-based methods. Specifically, for the first category of methods, the model’s decoder primarily utilizes the architecture of the ${{\rm TransEncoder}}$. As shown in Figure 6 (b), text tokens $T$ and video tokens $F_{v}$ are concatenated along the sequence dimension. Each of these tokens serves as the query, key, and value, which are fed into the transformer encoder to generate target text:
$T_{gen}=\{x_{t}\}_{t=1}^{N}={{\rm TransEncoder}}(Concat(T,F_{v})),$ (5)
$T_{gen}$ denotes the generated caption sequence of length $N$. For the second category of methods, the model’s decoder mainly adopts the structure of the ${{\rm TransDecoder}}$. Text tokens $T$ serves as the query, with video tokens $F_{v}$ acting as keys and values used for generating text:
$T_{gen}=\{x_{t}\}_{t=1}^{N}={{\rm TransDecoder}}(T,F_{v},F_{v}).$ (6)
For the above transformer-based methods, we concatenate subject tokens with video tokens along the sequence dimension. These models can utilize their attention mechanism to automatically learn representations to facilitate the generation of desired captions.
6 Experiments
6.1 Implementation Details
We employ four widely used metrics to assess the quality of the generated captions: BLEU@4 (B@4) [15], METEOR (M) [1], ROUGE-L (R) [8], and CIDEr (C) [23]. Bleu@4 focuses on how closely a translation matches the reference text, Rouge measures the completeness of the content in the generated text, Meteor values adaptability in different expressions, and CIDEr evaluates how the text relates to images. Thus, CIDEr serves as our primary evaluation metric.
We adopt the official codes of the four state-of-the-art methods and re-train them on our dataset using default hyperparameters considering they have similar data sizes. As for the object encoder employed in our new approach, we experimented with both Faster-RCNN [18] and CLIP [17]. The size of object features is 2048 obtained by Faster-RCNN. In the case of the pre-trained CLIP model, we fed the cropped regions directly into the image encoder with a VIT-L/14 backbone, leading to a feature dimension of 768.
6.2 Quantitative Results
Baselines Performance
In Table 2, we present the performance of the four baseline models and our extended models based on them. For a clear comparison, we also included their performance on the original MSVD and MSRVTT datasets, which are presented in rows with grey backgrounds.
The results show that: (I) The four state-of-the-art methods encounter a large performance drop on our new task compared to the performance on the original video captioning task. For example, on our subject-oriented MSVD, the CIDEr score of AR-B decreases by 31.7% relative to itself, while HMN reduces by 33.8% compared to its own performance. These indicate our subject-oriented video captioning task poses new challenges to existing methods. To succeed, they should have the ability to be aware of the specified target, and the ability to distinguish related activities. (II) Our extension of the four state-of-the-art methods brings large improvement. For example, on subject-oriented MSRVTT, our extended SwinBERT’s CIDEr score improves by 25.3% compared to without extension, while the extension of GL-RG exhibits an increase of 33.7%. These demonstrate the effectiveness of our utilization of user-specified targets. As these are intuitive extensions, we believe future dedicatedly designed methods could further boost the caption quality. (III) The performance of our extended methods still fall far behind their performance on the original video captioning tasks. For example, our best-performing extended SwinBERT on the subject-oriented MSVD, scored lower across all four performance metrics compared to the original SwinBERT in the conventional video captioning task, with a notable 35.2% decrease in BLEU@4. These indicate there is a large room for future exploration in this field.
Subject Encoder SO-MSVD SO-MSRVTT
B@4 M R C B@4 M R C
AR-B (Faster-RCNN) 30.3 23.8 52.5 76.2 19.3 17.6 41.8 50.8
AR-B (CLIP V/L-14) 33.5 25.3 53.8 86.9 19.6 17.8 40.6 50.2
HMN (Faster-RCNN) 32.7 24.1 53.6 85.2 19.8 17.6 42.5 52.2
HMN (CLIP V/L-14) 34.3 24.8 54.3 88.1 20.2 17.9 42.8 53.1
GL-RG (Faster-RCNN) 37.8 27.5 56.5 95.9 20.8 18.2 42.0 53.3
GL-RG (CLIP V/L-14) 37.2 27.8 56.9 97.6 22.2 18.9 43.9 58.6
SwinBert (Faster-RCNN) 35.6 25.9 55.9 98.1 21.3 19.1 43.9 58.0
SwinBert (CLIP V/L-14) 36.1 27.2 56.3 98.9 21.5 19.1 44.0 58.8
Table 3: Results under different visual subject encoders.
Effect of Different Subject Feature
In the computer vision community, there are two popular object encoders: Faster-RCNN and VIT-CLIP. To investigate the effect of different subject features on performance, we conduct tests using Faster-RCNN and VIT(L-14)-based CLIP to extract object features, respectively. The results are presented in Table 3. It shows that CLIP V/L-14 leads to better results for all methods on the SO-MSVD dataset in terms of the main metric Cider. The improvements range from 0.8 to 10.7. On the SO-MSRVTT dataset, the performance of three out of four methods is improved. Overall, using CLIP-VIT features achieves better performance.
Effect of Different Manners of Utilizing Subject Features
In addition to concatenating subject features with video features, we have also explored using addition to impose the visual subjects on video features. The results are present in Table 4. As shown in Table 4, for methods based on LSTM (i.e., HMN and GL-RG), using concatenation significantly respectively enhances CIDEr performance 4.5% and 19.0% on the SO-MSVD dataset. In contrast, for Transformer-based methods, the difference between concatenation and addition is marginal.
The main reason can be attributed to the fact that in LSTM-based methods, the length of subject features has to be the same as the video feature for the addition operation, while it becomes more flexible in the concatenation operation. In contrast, in Transformer-based methods, as the concatenation is performed along the sequence dimension, the feature dimension has to be the same as other tokens.
Combination Operations SO-MSVD SO-MSRVTT
B@4 M R C B@4 M R C
AR-B (Add) 33.0 25.1 53.6 88.2 19.4 17.7 40.5 49.6
AR-B (Concat) 33.5 25.3 53.8 86.9 19.6 17.8 40.6 50.2
HMN (Add) 31.9 23.9 53.0 84.3 19.3 17.3 41.2 50.3
HMN (Concat) 34.3 24.8 54.3 88.1 20.2 17.9 42.8 53.1
GL-RG (Add) 35.9 25.5 54.5 82.0 22.1 18.9 43.6 55.9
GL-RG (Concat) 37.2 27.8 56.9 97.6 22.2 18.9 43.9 58.6
SwinBert (Add) 36.6 27.2 56.5 97.2 21.7 18.9 44.5 59.0
SwinBert (Concat) 36.1 27.2 56.3 98.9 21.5 19.1 44.0 58.8
Table 4: Results under different combination operations based on CLIP subject features.
Effect of Different Dimension of Subject Features
Figure 7: Results (CIEDr score) on Subject-Oriented MSVD and MSRVTT under different subject feature dimensions.
For LSTM-based methods, the dimensionality of the subject features can significantly impact the performance of the generated captions, as illustrated in Figure 7. For SO-MSVD, lower dimensions of subject features tend to perform better, while for SO-MSRVTT, higher dimensions show improved performance. This is because the SO-MSRVTT dataset is larger and more complex, and higher feature dimensions help the model to better fit and capture the diversity and complexity of subject features.
Figure 8: Qualitative results on Subject-Oriented MSR-VTT.
Accuracy of Predicted Subject
In addition to using four conventional video description quality assessment metrics, we also compared the subject accuracy of the modified methods against the original methods on the test set of the new datasets. If the predicted subject is found within the set of ground truth subjects, the prediction is considered correct; otherwise, it is deemed incorrect.
The results, as shown in Table 5, indicate that our method achieves improvements over the baseline. Specifically, on the SO-MSRVTT dataset, SwinBERT and GL-RG achieved significant enhancements of approximately 12.8% and 11.8%, respectively. Meanwhile, on the SO-MSVD dataset, AR-B showed an improvement of around 9.8%.
Methods SO-MSVD SO-MSRVTT
Acc Acc
AR-B 52.5% 46.5%
AR-B (Ours) 62.3% 50.1%
HMN 59.6% 47.4%
HMN (Ours) 65.2% 60.8%
GL-RG 56.5% 45.6%
GL-RG (Ours) 60.0% 57.4%
SwinBert 60.7% 48.8%
SwinBert (Ours) 63.9% 61.6%
Table 5: Subject accuracy of different methods on our Subject-Oriented MSVD and MSR-VTT benchmarks.
6.3 Qualitative Results
We show several qualitative results on our subject-oriented MSRVTT in Figure 8. Our improved method, compared to the four models that do not utilize subject areas, better focuses on and accurately describes the entities of our interest. For instance, the video in Figure 8 (a) shows three people conversing. It can be observed that the four original methods only described two of them. In contrast, our method generates accurate descriptions for the individually marked single, double, and triple-person groups in the video. Figure 8 (b) displays a group of people visiting a large boat. SwinBERT incorrectly identifies the boat in the video as “A large train”, while the other three methods describe the people. Our method can accurately describe both “A boat is being shown” and “a man is watching a boat”. In Figure 8 (c), the basic methods only describe people. However, our method not only describes “A woman is talking” and “A couple of people are walking down the street”, but also notes “A white building is being filmed”. In Figure 8 (d), our method is capable of describing “A machine” appearing in the video, an entity not mentioned by the original methods. The qualitative results further prove the effectiveness of our method and the significance of this task.
6.4 Future Work
The task of subject-oriented video caption generation is highly challenging. It requires the model to not only understand the video content but also to focus specifically on the behavior of entities that are of interest to the user. How to more effectively utilize the subject area is one of the possible research directions for this task. Future work may also explore large-scale vision-language models, such as GPT-4V.
7 Conclusion
In this paper, we introduce a new task, Subject-Oriented video captioning, which aims at describing the activities of entities in videos that are of interest to users. We build two datasets to support future research in this field and evaluate four state-of-the-art conventional video captioning methods as well as our extensions based on them. Experimental results show a large performance drop of existing methods on our new tasks and significant improvement of our extended models. Our extended methods can serve as strong baselines, but there is still a large room worth exploring for future works.